The camera is hung in a state that requires a power cycling.
Camera fails to execute a software reset
The camera returns the error “failure in reset”
The leach controller window shows highly incorrect values in gain (e.g., -4095), amp, heat, etc.
Cycle the power on the camera on site (cannot be done remotely). It is unclear where this problem stems from.
Cloud limit was raised to 70%. This should allow some backlogged requests to get through.
There was an unknown error with the leach controller daemon last night. The daemon was restarted, and now appears to be working.
Please ignore the nomets messages in last night’s queue. We were running a test. Weather was to poor to observe.
The MAWS is back and working fine. The serial port the MAWS was connected to showed signs of intermittent failure. A backup card was installed, and the port changed (ttyS0 -> ttyS4). The opportunity was taken to re-evacuate the dewar, which was not cooling completely. The telescope is on and functioning.
The vaisala weather station is not responding. The telescope will be down for at least a few days until the station can be inspected.
I’ve updated the RCT webpage to highlight science. I’ve moved the telescope maintenance issues and bugs to a separate location on the site, but you will still get posts will still get to the RSS. Please review and let me know if have suggestions or comments.
The chip orientation has been erroneous since moving to amp=C a long while ago. The chip orientation should now be correct with North=Left East=Down.
The focus was off last night. I’ll interrupt the queue at the beginning of tonight to re-calibrate it.